George and elena made $58 selling candy. George made $6 more than 3 times as much money as elena made. How much money did George and elena make?
Answer by TimothyLamb(4379)

x = dollars george made
y = dollars elena made
---
x + y = 58
x = 3y + 6
---
put the system of linear equations into standard form
---
x + y = 58
x - 3y = 6
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = dollars george made = 45
y = dollars elena made = 13
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---